OpenStudy (anonymous):

How do i do X^8 -X^5 i tried taking out a gcf of X^5 but i dont know if thats right ? If so where to go from their or whats the correct way ? thanks :D

OpenStudy (cwrw238):

x^5 is the correct gcf now divide x^8 and x^5 by x^5 x^8 / x^5 = x^(8-5) = x^3 and x^5 / x^5 is of course 1

OpenStudy (cwrw238):

so the factors are x^5 (x^3 - 1)

OpenStudy (phi):

you could go a little further: x^3-1= (x-1)(x^2+x+1) so you could write it as \[x^5(x^2+x+1)(x-1)\]

OpenStudy (anonymous):

Thanks so much to the two of you ! :D